# Earliest Galaxy Evolution in the CANUCS+Technicolor fields: Galaxy Properties at $z\sim 10-16$ seen with the Full NIRCam Medium and Broad Band Filters

### v1

@@ 3.1 A remarkable $z>15$ galaxy candidate: CANUCS-3214552
The highest redshift candidate among the sample is CANUCS-3214552, whose best photo-$z$ estimate is $z_{\rm ml}=15.43^{+0.26}_{-0.57}$ with a negligible low-$z$ possibility ($P(z<7)=5.1\times 10^{-3}$; Figure 1). The galaxy is faint and relatively red in UV ($M_{\rm UV}=-19.0$ mag and $\beta_{\rm UV}=-1.8$), and the sharp dropout is captured between the F182M+F200W and F210M filters. Together with the flat SED at longer wavelengths, the sharp break in the narrow wavelength range between F182M and F210M and the dropout in F200W strongly support the $z=15.4$ solution over the low-$z$ ($z=4$) solution (the blue solid curve in the bottom right panel of Figure 1). Indeed, this source would have a non-negligible secondary low-$z$ peak in the PDZ if only the standard set of NIRCam eight filters (seven BBs + F410M) were available (the dashed curve in the panel) and would not be selected as a secure high-$z$ candidate.
Figure 1: The highest redshift candidate of CANUCS-3214552 at $z_{\rm phot}\sim 15.4$. Top: cutouts of the source in all NIRCam filter images, except for the two narrow band filters. Cutouts are $1.\!\!^{\prime\prime}6$ on the side. Bottom: the SED and the PDZ of CANUCS-3214552. The SED shows a clear dropout between F210M and F200W, and the sharp drop between this narrow wavelength range strongly supports the high-$z$ solution template (black solid curve in the bottom left panel) and results in negligible low-$z$ probability (blue curve in the bottom right panel). There is no variability between Cycle 1 and Cycle 2 observation with 1 year separation (blue and red squares in bottom left), which rules out the possibility of $z\sim 4$ SN contamination. If MB data were not available, the SED could also be fit by a low-$z$ solution (gray curve in bottom left) and the PDZ would have a non-negligible secondary peak at $z\sim 4$ (gray dashed line in bottom right). Only multiple-epoch rich MB+BB observations can secure this galaxy as a very robust $z\sim 15$ galaxy candidate.
Moreover, the multi-epoch observations confirmed there is no evidence for variability. Several studies have shown that some types of supernovae (SNe) at $52<z<$ 5 have strong breaks around rest 4000 Å wavelength, which can mimic the color of highest-$z$ LBGs (e.g., H. Yan et al., 2023; C. DeCoursey et al., 2025). C. DeCoursey et al. (2025) suggested that a high fraction of point-like, $z\sim 16$ candidates could be SNe at $z\sim 4$. However in the case of CANUCS-3214552, the galaxy was observed in Cycle 1 and Cycle 2 with one year separation and the NIRCam photometry agrees well between the two epochs (e.g., F335M and F360M taken in Cycle 1 vs. F356W taken in Cycle 2).
CANUCS-3214552 is thus one of the most reliable high-$z$ galaxy candidates at $z>15$, which could have been secured only with the multi-epoch, full NIRCam MB+BB observations. The galaxy shows a relatively red rest-UV color, in contrast to recent reports of very blue colors for the most of faint highest-$z$ galaxies (e.g., F. Cullen et al., 2024), and it could suggest the galaxy represents a galaxy population that has been missed in previous surveys using the standard NIRCam BB configuration. We provide a detailed discussion on this point later (Sec. 5.3.2).
@@ 3.2 Completeness simulation
A luminosity function estimation requires computation of the effective volume that has been searched, taking account of the effect of detection and selection completeness. Our selection criteria (equation 1) are fully automated and we do not apply additional cuts based on visual inspection. Thus, we can estimate the completeness by inserting simulated sources into our data and repeating the detection and selection procedure to measure the recovered fraction. We follow the completeness simulation prescription of C. J. Willott et al. (2024) with a few modifications.
For the simulated galaxy size distribution, C. J. Willott et al. (2024) used a redshift-dependent size-$M_{\rm UV}$ relationship derived from a photometric-redshift-selected sample of high-$z$ galaxies from T. Morishita et al. (2024). As we show in Sec. 5.3.1, the $z>9.5$ galaxies from the current study plus those from the JOF that also includes MB photometry (B. Robertson et al., 2024) have sizes considerably smaller (median offset 0.6 dex) than that of the T. Morishita et al. (2024) relationship. Therefore, we offset the T. Morishita et al. (2024) relationship by 0.6 dex for the size distribution of the simulated galaxies.
@@
A grid of $M_{\rm UV}$ and redshift covering the range of this study was set up. At each point in this plane, simulated galaxies were randomly inserted into all the NIRCam images of the three fields. In total, 850,000 galaxies were inserted, but not more than 500 per NIRCam module at a time to prevent crowding of simulated sources. These images were then analyzed using the same detection and segmentation process, followed by the sample selection criteria of equation 1. As in C. J. Willott et al. (2024), it is found that $\sim 10\%$ of bright objects are not recovered due to overlap with existing galaxies in the images.
Figure 2 shows the completeness of our high-$z$ galaxy selection in the $M_{\rm UV}$ vs. redshift plane. Our selection is $\sim 50\ \%$ complete down to $M_{\rm UV}\sim-18.7$ at $z\gtrsim 10$. The observed $M_{\rm UV}$ and redshift values of our high-$z$ sample are also plotted in the figure (yellow points). The distribution of observed galaxies agrees well with the predicted completeness from the simulation (e.g., only a few galaxies in the $<50\ \%$ complete area), which supports the validity of our simulation configuration. In particular, if we were to use the original size-$M_{\rm UV}$ relation of T. Morishita et al. (2024) in the simulation, five of the eight galaxies in our sample would be located in the $<50\ \%$ completeness region in the $M_{\rm UV}$-$z$ plane.
@@ 4.1 Evolving UV Luminosity Functions
Knowing the completeness of our survey and high-$z$ galaxy selection, we can compute the UV luminosity function and its redshift evolution at $z\gtrsim 10$. Although our selection is nearly complete down to $M_{\rm UV}\sim-19$ mag beyond $z=12$, there is only one galaxy found in the total $\sim 23\ {\rm arcmin}^{2}$ unmasked area at $z>12$, which suggests a significant evolution of the galaxy UV luminosity function at these redshifts. We parameterize the UV LF in an analytical form with redshift-dependent parameters, and directly fit to data without binning in redshift or magnitude to avoid any effect by binning, considering the paucity of $z\gtrsim 11$ galaxies. We adopt the Schechter form (P. Schechter, 1976) for the LF, {dmath} ϕ_UV(M_UV,z) = ln(10)2.5ϕ_⋆(z) 10^0.4(α+1)(M_UV^⋆-M_UV) exp[-10^0.4(M_UV^⋆-M_UV)], where $\alpha$ and $M_{\rm UV}^{\star}$ are the faint-end slope and the characteristic UV absolute magnitude, respectively, and $\phi_{\star}(z)$ is the redshift-dependent normalization parameter. We assume a simple redshift dependency, as done by B. Robertson et al. (2024),
@@
to reduce the free parameters in the fit. The reference redshift $z_{0}$ is fixed to $z_{0}=10$ throughout the paper unless otherwise specified. We also fix $\alpha$ and $M_{\rm UV}^{\star}$ to $\alpha=-2.1$ and $M_{\rm UV}^{\star}=-21.0$ mag referring to previous studies of UV LFs at $z>10$ (e.g., N. J. Adams et al., 2024; C. T. Donnan et al., 2024; B. Robertson et al., 2024; C. J. Willott et al., 2024). Thus the model LF has two free parameters of $\phi_{\star,0}$ and $\eta$.
We follow H. L. Marshall et al. (1983) for the LF parameter fitting, so that we can also take into account the non-detections in the relatively bright (brighter than $M_{\rm UV}\sim-19$ mag) galaxies at $z\gtrsim 12$. We derive the best estimations of $\phi_{\star,0}$ and $\eta$ to minimize the function $S=-2\ln\mathcal{L}$, where $\mathcal{L}$ is the likelihood and {dmath} lnL = ∑_i^N [ ϕ_UV(M_UV i, z_i) p(M_UV i, z_i) ] - ∬ϕ_UV(M_UV,z)p(M_UV, z) dVdz dz dM_UV. The first term is summed over all observed high-$z$ sample galaxies $i$, and the second term is the integral over all possible range of $M_{\rm UV}$ and $z$. $p(M_{\rm UV},z)$ is the completeness at the point of $(M_{\rm UV},z)$, as estimated in Sec. 3.2, and $dV/dz$ is the differential volume element at $z$. The combination of $p(M_{\rm UV},z)$ and $dV/dz$ can thus consider the effective sky area of all three fields at $(M_{\rm UV},z)$, taking account for the completeness. We use a python implementation of the MCMC (emcee; D. Foreman-Mackey et al., 2013) to obtain the posterior distributions of $\phi_{\star,0}$ and $\eta$, assuming flat priors between $-8<\log_{10}\phi_{\star,0}<-2$ and $-3<\eta<3$. Table 3 present the fitting results. The best estimations are from the median of the marginalized posterior distribution, and the uncertainties are from 16th- and 84th-percentiles of the posterior. We here stress again that the fit is based on all redshift ranges where the completeness has non-negligible value and also accounts for the non-detections of galaxies in the $M_{\rm UV}$-$z$ plane.
Table 3: Parameters of redshift-dependent Schechter UV LFs at $z>9.5$ inferred from CANUCS+TEC data.
@@
$\log_{10}(\phi_{\star,0}/{\rm Mpc^{-3}\ mag^{-1}})$ $\mathcal{U}(-8,-2)$ $-4.89^{+0.17}_{-0.20}$
$\eta$ $\mathcal{U}(-3,3)$ $-0.21^{+0.11}_{-0.12}$
$\alpha$ $-2.1^{*}$
$M_{\rm UV}^{\star}/{\rm mag}$ $-21.0^{*}$
footnotetext: Notes. The normalization parameter $\phi_{\star}$ is assumed to evolve depending on redshifts as $\log_{10}\phi_{\star}(z)=\log_{10}(\phi_{\star,0})+\eta(z-10)$, and the fitting results for the two parameters $\eta$ and $\log_{10}(\phi_{\star,0})$ are given. Values with an asterisk are fixed in the fit.
The log-linear rate of the UV LF redshift evolution $\eta$ obtained in this work ($\eta=-0.21^{+0.11}_{-0.12}$) is consistent with previous work in the JOF that used a similarly rich NIRCam MB+BB filter set (B. Robertson et al., 2024). This means that the logarithmic galaxy abundance declines at a rate of $-0.21$ per unit redshift at $z\gtrsim 10$; for example, the number density should decrease by a factor of $\sim 7$ from $z\sim 11$ to $z\sim 15$. On the other hand, the normalization parameter at the reference redshift $\phi_{\star,0}$ in this work is smaller than B. Robertson et al. (2024) by a factor of $\sim 0.6$ dex (after correcting the difference in the reference redshift $z_{0}$ and degeneracy between $\phi_{\star,0}$ and $M_{\rm UV}^{\star}$). This implies that cosmic variance (CV) can significantly affect galaxy number density measurement by at least $0.6$ dex.
Figure 3: The UV LFs at $z\sim 11$ (left) and $z\sim 14$ (right) from CANUCS+TEC data. Black solid curve is the best estimation of the redshift-evolving UV LF obtained in Sec. 4.1 at $z=10.5$ (left) and $z=14$ (right), with the red-shaded area showing 16th- to 84th-percentile. Red filled circles with black error bars present the binned UV LFs in the $129.5<z<$ 12 bin (left) and $1612<z<$ 16 bin (right) measured in Sec. 4.2. UV LF measurements in literature at similar redshifts are also shown for comparison (P. G. Pérez-González et al., 2023; N. J. Adams et al., 2024; C. T. Donnan et al., 2024; B. Robertson et al., 2024; C. J. Willott et al., 2024; P. G. Pérez-González et al., 2025; M. Castellano et al., 2025).
@@ 4.2 Binned UV LFs
Although we estimate the evolving UV LFs without binning the sample galaxies in the previous section, deriving the binned LFs is still useful to compare with LF measurements in literature. We use the binned $1/V_{a}$ method of Y. Avni & J. N. Bahcall (1980) and obtain the LFs at $129.5<z<$ 12 and $1612<z<$ 16 separately. The available effective volume for a galaxy $i$ in a redshift/magnitude bin with the size of $\Delta z$ and $\Delta M_{\rm UV}$ is
[…]
@@
Uncertainties on the number density is based on Poisson statistics with the correction for small numbers by N. Gehrels (1986).
Table 4: Binned UV LFs
@@
footnotetext: Notes. Upper limits are $1\sigma$.
Table 4 and Figure 3 presents the result. We quote the $1\sigma$ upper limits when no galaxy is detected in a bin. In the figure, we also show the marginalized LFs at $z=10.5$ and $z=14$ obtained from the MCMC sampling in Sec. 4.1 (black solid curves with red shaded regions). In the lower redshift bin at $z\sim 11$ (left panel in Figure 3), our LF measurements are consistent with or at the lower envelope of other recent JWST studies (P. G. Pérez-González et al., 2023; N. J. Adams et al., 2024; C. T. Donnan et al., 2024; B. Robertson et al., 2024; C. J. Willott et al., 2024; L. Whitler et al., 2025). The spread of LF measurements across the literature is roughly $\sim 0.6$ dex, which can be understood as the effect of the CV. In the higher redshift bin at $z\sim 14$ (right panel in the figure), although we have only one detection in the faintest magnitude bin, the non-detection of relatively bright ($M_{\rm UV}<-19$ mag) galaxies in the fairly large survey volume puts a tight constraint on the UV LFs, which results in a decline of galaxy abundance at $z>12$ ($\eta=-0.2$ as obtained in Sec.4.1). We note that our LF measurement at $z\sim 14$ is comparable to or even lower than recent reports of UV LFs at $z\sim 17$ (M. Castellano et al., 2025; P. G. Pérez-González et al., 2025), which mostly used only NIRCam BB data to select high-$z$ galaxy candidates. This is possibly due to the low-$z$ interloper contaminants when selecting high-$z$ candidates only based on BB-only catalogs, rather than the effect of CV or very shallow UV LF evolution at $z>15$ (see Sec. 5.2).

### v2

@@ 3.1 A remarkable $z>15$ galaxy candidate: CANUCS-3214552
[…]
Figure 2: The highest redshift candidate of CANUCS-3214552 at $z_{\rm phot}\sim 15.4$. Top: cutouts of the source in all NIRCam filter images, except for the two narrow band filters. Cutouts are $1.\!\!^{\prime\prime}6$ on the side. Image scalings are homogeneous. Bottom: the SED and the PDZ of CANUCS-3214552. The SED shows a clear dropout between F210M and F200W, and the sharp drop between this narrow wavelength range strongly supports the high-$z$ solution template (black solid curve in the bottom left panel) and results in negligible low-$z$ probability (blue curve in the bottom right panel). There is no variability between Cycle 1 and Cycle 2 observation with 1 year separation (blue and red squares in bottom left), which rules out the possibility of $z\sim 4$ SN contamination. If MB data were not available, the SED could also be fit by a low-$z$ solution (gray curve in bottom left) and the PDZ would have a non-negligible secondary peak at $z\sim 4$ (gray dashed line in bottom right). Only multiple-epoch rich MB+BB observations can secure this galaxy as a very robust $z\sim 15$ galaxy candidate.
Moreover, the multi-epoch observations confirmed there is no evidence for variability. Several studies have shown that some types of supernovae (SNe) at $2<z<5$ have strong breaks around rest 4000 Å wavelength, which can mimic the color of highest-$z$ LBGs (e.g., Yan2023ApJS; DeCoursey2025ApJ). DeCoursey2025ApJ suggested that a high fraction of point-like, $z\sim 16$ candidates could be SNe at $z\sim 4$. However in the case of CANUCS-3214552, the galaxy was observed in Cycle 1 and Cycle 2 with one year separation and the NIRCam photometry agrees well between the two epochs (e.g., F335M and F360M taken in Cycle 1 vs. F356W taken in Cycle 2).
CANUCS-3214552 is thus one of the most reliable high-$z$ galaxy candidates at $z>15$, which could have been secured only with the multi-epoch, full NIRCam MB+BB observations. The galaxy shows a relatively red rest-UV color, in contrast to recent reports of very blue colors for the most of faint highest-$z$ galaxies (e.g., Cullen2024MNRAS), and it could suggest the galaxy represents a galaxy population that has been missed in previous surveys using the standard NIRCam BB configuration. We provide a detailed discussion on this point later (Sec. 5.3.2).
@@ 3.2 Completeness simulation
A luminosity function estimation requires computation of the effective volume that has been searched, taking account of the effect of detection and selection completeness. Our selection criteria (equation 1) are fully automated and we do not apply additional cuts based on visual inspection. Thus, we can estimate the completeness by inserting simulated sources into our data and repeating the detection and selection procedure to measure the recovered fraction. We follow the completeness simulation prescription of Willott2024ApJ with a few modifications.
For the simulated galaxy size distribution, Willott2024ApJ used a redshift-dependent size-$M_{\rm UV}$ relationship derived from a photometric-redshift-selected sample of high-$z$ galaxies from Morishita2024. As we show in Sec. 5.3.1, the $z>9.5$ galaxies from the current study plus those from the JOF that also includes MB photometry (Robertson2024ApJ) have sizes considerably smaller (median offset 0.6 dex) than that of the Morishita2024 relationship. Therefore, we offset the Morishita2024 relationship by 0.6 dex for the size distribution of the simulated galaxies.
@@
A grid of $M_{\rm UV}$ and redshift covering the range of this study was set up. At each point in this plane, simulated galaxies were randomly inserted into all the NIRCam images of the three fields. In total, 850,000 galaxies were inserted, but not more than 500 per NIRCam module at a time to prevent crowding of simulated sources. These images were then analyzed using the same detection and segmentation process, followed by the sample selection criteria of equation 1. As in Willott2024ApJ, it is found that $\sim 10\%$ of bright objects are not recovered due to overlap with existing galaxies in the images.
Figure 3 shows the completeness of our high-$z$ galaxy selection in the $M_{\rm UV}$ vs. redshift plane. Our selection is $\sim 50\ \%$ complete down to $M_{\rm UV}\sim-18.7$ at $z\gtrsim 10$. The observed $M_{\rm UV}$ and redshift values of our high-$z$ sample are also plotted in the figure (yellow points). The distribution of observed galaxies agrees well with the predicted completeness from the simulation (e.g., only a few galaxies in the <50%<50\ $\%$ complete area), which supports the validity of our simulation configuration. In particular, if we were to use the original size-$M_{\rm UV}$ relation of Morishita2024 in the simulation, five of the eight galaxies in our sample would be located in the <50%<50\ $\%$ completeness region in the $M_{\rm UV}$-$z$ plane.
@@ 4.1 Evolving UV Luminosity Functions
Knowing the completeness of our survey and high-$z$ galaxy selection, we can compute the UV luminosity function and its redshift evolution at $z\gtrsim 10$. Although our selection is nearly complete down to $M_{\rm UV}\sim-19$ mag beyond $z=12$, there is only one galaxy found in the total $\sim 23\ {\rm arcmin}^{2}$ unmasked area at $z>12$, which suggests a significant evolution of the galaxy UV luminosity function at these redshifts. We parameterize the UV LF in an analytical form with redshift-dependent parameters, and directly fit to data without binning in redshift or magnitude to avoid any effect by binning, considering the paucity of $z\gtrsim 11$ galaxies. We adopt the Schechter form (Schechter1976ApJ) for the LF,
$\phi_{\rm UV}(M_{\rm UV},z)=\frac{\ln(10)}{2.5}\phi_{\star}(z)\ 10^{0.4(\alpha+1)(M_{\rm UV}^{\star}-M_{\rm UV})}\exp\left[-10^{0.4(M_{\rm UV}^{\star}-M_{\rm UV})}\right],$ (2)
where $\alpha$ and $M_{\rm UV}^{\star}$ are the faint-end slope and the characteristic UV absolute magnitude, respectively, and $\phi_{\star}(z)$ is the redshift-dependent normalization parameter. We assume a simple redshift dependency, as done by Robertson2024ApJ,
@@
to reduce the free parameters in the fit. The reference redshift $z_{0}$ is fixed to $z_{0}=10$ throughout the paper unless otherwise specified. We also fix $\alpha$ and $M_{\rm UV}^{\star}$ to $\alpha=-2.1$ and $M_{\rm UV}^{\star}=-21.0$ mag referring to previous studies of UV LFs at $z>10$ (e.g., Adams2024ApJ; Donnan2024; Robertson2024ApJ; Willott2024ApJ). Thus the model LF has two free parameters of $\phi_{\star,0}$ and $\eta$. Although there are some evidence preferring the double-power law function over the Schechter function in high-$z$ UV LFs (e.g., Harikane2022ApJS; Donnan2024), we assume the Schechter function in this work because we do not have a constraint on the bright-end of LFs to distinguish the Schechter function and double-power law and the Schechter function has one fewer parameter. We confirmed that all results in this paper do not change even if we adopt the double-power law function instead of the Schechter form.
We follow Marshall1983 for the LF parameter fitting, so that we can also take into account the non-detections in the relatively bright (brighter than $M_{\rm UV}\sim-19$ mag) galaxies at $z\gtrsim 12$. We derive the best estimations of $\phi_{\star,0}$ and $\eta$ to minimize the function $S=-2\ln\mathcal{L}$, where $\mathcal{L}$ is the likelihood and
$\ln\mathcal{L}=\sum_{i}^{N}\left[\phi_{\rm UV}(M_{{\rm UV}\ i},z_{i})p(M_{{\rm UV}\ i},z_{i})\right]-\iint\phi_{\rm UV}(M_{\rm UV},z)p(M_{\rm UV},z)\frac{dV}{dz}\ dz\ dM_{\rm UV}.$ (4)
The first term is summed over all observed high-$z$ sample galaxies $i$, and the second term is the integral over all possible range of $M_{\rm UV}$ and $z$. $p(M_{\rm UV},z)$ is the completeness at the point of $(M_{\rm UV},z)$, as estimated in Sec. 3.2, and $dV/dz$ is the differential volume element at $z$. The combination of $p(M_{\rm UV},z)$ and $dV/dz$ can thus consider the effective sky area of all three fields at $(M_{\rm UV},z)$, taking account for the completeness. We use a python implementation of the MCMC (emcee; emcee) to obtain the posterior distributions of $\phi_{\star,0}$ and $\eta$, assuming flat priors between $8<\log_{10}\phi_{\star,0}<-2$ and $3<\eta<3$. Table 3 present the fitting results. The best estimations are from the median of the marginalized posterior distribution, and the uncertainties are from 16th- and 84th-percentiles of the posterior. We here stress again that the fit is based on all redshift ranges where the completeness has non-negligible value and also accounts for the non-detections of galaxies in the $M_{\rm UV}$-$z$ plane.
Table 3: Parameters of redshift-dependent Schechter UV LFs at $z>9.5$ inferred from CANUCS+TEC data. footnotetext: Notes. The normalization parameter $\phi_{\star}$ is assumed to evolve depending on redshifts as $\log_{10}\phi_{\star}(z)=\log_{10}(\phi_{\star,0})+\eta(z-10)$, and the fitting results for the two parameters $\eta$ and $\log_{10}(\phi_{\star,0})$ are given. Values with an asterisk are fixed in the fit.
@@
$\log_{10}(\phi_{\star,0}/{\rm Mpc^{-3}\ mag^{-1}})$ $\mathcal{U}(-8,-2)$ $4.89^{+0.17}_{-0.20}$
$\eta$ $\mathcal{U}(-3,3)$ $0.21^{+0.11}_{-0.12}$
$\alpha$ $2.1^{*}$
$M_{\rm UV}^{\star}/{\rm mag}$ $21.0^{*}$
The log-linear rate of the UV LF redshift evolution $\eta$ obtained in this work ($\eta=-0.21^{+0.11}_{-0.12}$) is consistent with previous work in the JOF that used a similarly rich NIRCam MB+BB filter set (Robertson2024ApJ). This means that the logarithmic galaxy abundance declines at a rate of 0.21-0.21 per unit redshift at $z\gtrsim 10$; for example, the number density should decrease by a factor of $\sim 7$ from $z\sim 11$ to $z\sim 15$. On the other hand, the normalization parameter at the reference redshift $\phi_{\star,0}$ in this work is smaller than Robertson2024ApJ by a factor of $\sim 0.6$ dex (after correcting the difference in the reference redshift $z_{0}$ and degeneracy between $\phi_{\star,0}$ and $M_{\rm UV}^{\star}$). This implies that cosmic variance (CV) can significantly affect galaxy number density measurement by at least $0.6$ dex.
Figure 4: The UV LFs at $z\sim 11$ (left) and $z\sim 14$ (right) from CANUCS+TEC data. Black solid curve is the best estimation of the redshift-evolving UV LF obtained in Sec. 4.1 at $z=10.5$ (left) and $z=14$ (right), with the red-shaded area showing 16th- to 84th-percentile. Red filled circles with black error bars present the binned UV LFs in the $9.5<z<12$ bin (left) and $12<z<16$ bin (right) measured in Sec. 4.2. UV LF measurements in literature at similar redshifts are also shown for comparison (Perez_Gonzalez2023; Adams2024ApJ; Donnan2024; Robertson2024ApJ; Willott2024ApJ; Perez_Gonzalez2025; Castellano2025arXiv).
@@ 4.2 Binned UV LFs
Although we estimate the evolving UV LFs without binning the sample galaxies in the previous section, deriving the binned LFs is still useful to compare with LF measurements in literature. We use the binned $1/V_{a}$ method of Avni1980 and obtain the LFs at $9.5<z<12$ and $12<z<16$ separately. The available effective volume for a galaxy $i$ in a redshift/magnitude bin with the size of $\Delta z$ and $\Delta M_{\rm UV}$ is
$V_{a}^{i}=\iint p(M_{\rm UV},z)\frac{dV}{dz}\ dz\ dM_{\rm UV},$ (5)
@@
Uncertainties on the number density is based on Poisson statistics with the correction for small numbers by Gehrels1986.
Table 4: Binned UV LFs footnotetext: Notes. Upper limits are $1\sigma$.
@@
Table 4 and Figure 4 presents the result. We quote the $1\sigma$ upper limits when no galaxy is detected in a bin. In the figure, we also show the marginalized LFs at $z=10.5$ and $z=14$ obtained from the MCMC sampling in Sec. 4.1 (black solid curves with red shaded regions). In the lower redshift bin at $z\sim 11$ (left panel in Figure 4), our LF measurements are consistent with or at the lower envelope of other recent JWST studies (Perez_Gonzalez2023; Adams2024ApJ; Donnan2024; Robertson2024ApJ; Willott2024ApJ; Whitler2025arXiv). The spread of LF measurements across the literature is roughly $\sim 0.6$ dex, which can be understood as the effect of the CV. In the higher redshift bin at $z\sim 14$ (right panel in the figure), although we have only one detection in the faintest magnitude bin, the non-detection of relatively bright ($M_{\rm UV}<-19$ mag) galaxies in the fairly large survey volume puts a tight constraint on the UV LFs, which results in a decline of galaxy abundance at $z>12$ ($\eta=-0.2$ as obtained in Sec.4.1). We note that our LF measurement at $z\sim 14$ is comparable to or even lower than recent reports of UV LFs at $z\sim 17$ (Castellano2025arXiv; Perez_Gonzalez2025), which mostly used only NIRCam BB data to select high-$z$ galaxy candidates. This is possibly due to the low-$z$ interloper contaminants when selecting high-$z$ candidates only based on BB-only catalogs, rather than the effect of CV or very shallow UV LF evolution at $z>15$ (see Sec. 5.2).